julio is 4 times as old as maria his sister. in 6 years he will only be twice as old. how old is julio and maria now

Answer by gwendolyn(128)

We want to determine the ages of Julio and Maria, so we will assign them both variables:
let j= Julio's age
let m= Maria's age
Now, we can use these variables in equations. The problem states that Julio is four 4 times as old as Maria. This means that Julio's age is four multiplied by Maria's age. This can be set up as the equation below:
j=4m
We have two variables, so we need another equation. We also know that in six years, Julio's age will be twice (two times) Maria's age. The way to write this is a little bit more complicated. First, we have to remember that this equation is from "in 6 years." Because of this, we need to add 6 to each age. [(j+6); (m+6)] Since Julio's age six years from now will be twice Maria's age 6 years from now, we can write the equation as below:
j+6=(m+6)*2
Since we already know j's value in terms of m (from above, j = 4m), we can substitute 4m for the j in our second equation:
4m + 6 = (m + 6)*2
Both halves of the equation can be simplified.
First, we'll distribute the two in the second half of the equation:
4m + 6 = 2m + 2*6
Then we can multiply the 6 and 2 in the second half:
4m + 6 = 2m + 12
After that, we can subtract 6 from each side:
4m + 6 -6 = 2m + 12 - 6
4m = 2m + 6
Now, we can subtract 2m from both sides.
4m - 2m = 2m + 6 -2m
2m = 6
And divide each side by 2:
m =3
Therefore, Maria is 3 years old.
We also need to determine Julio's age. We know Julio is 4 times Maria's age so we can plug Maria's age into that equation:
j=4m
j = 4*3
j = 12
So, Julio is 12 years old and Maria is 3 years old.